OpenStudy (anonymous):

5/7x-3=2 solve for x.

OpenStudy (anonymous):

Is it \[\frac{ 5 }{ 7x }-3 = 2 ~~~\text{or}~~~\frac{ 5 }{ 7 }x-3 =2\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Alright, so any idea how to start?

OpenStudy (anonymous):

I'm assuming you add 3 to both sides, right?

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

So do that now, and then try thinking of what the next step is :)

OpenStudy (anonymous):

multiply the reciprocal?

OpenStudy (anonymous):

Yeah! \[\frac{ 5 }{ 7 } \times \frac{ 7 }{ 5 }\]

OpenStudy (anonymous):

But multiply by 7/5 on both sides :)